OpenStudy (anonymous):

If tanθ=− 5/12 with −π/2<θ<0 , find the exact value of the other 5 trig functions of θ . If tanθ=− 5/12 with −π/2<θ<0 , find the exact value of the other 5 trig functions of θ . @Mathematics

OpenStudy (mertsj):

Have I taught you nothing yet??? This angle is in the fourth quadrant. If the tangent is 5/12, the side opposite is 5, the side adjacent is 12 and the hypotenuse is 13. so... sin = -5/13 (4th quadrant, y is negative) cos = 12/13 (4th quadrant, x is positive) cot = -12/5 (reciprocal of tan) sec = 13/12 (reciprocal of cos) csc = -13/5 (reciprocal of sin)

OpenStudy (anonymous):

whenever i draw the triangle,where will i call my opposite?